UT 20150601 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson, BRothberg
Observer AO: DMiller
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Will loose some part of the first half of the night due to clouds. If it clears up as expected
will continue with some AO-stuff, ie image quality and PSF-tests

Summary

- AO image-quality and PSF-tests

- Lost some time at begin of night due to clouds

Details

All times are UT
Files are luci2.20150601.NNNN.fits

02:25 Sunset, mostly cloudy (fat cloud above us), scattered showers around,
T = +12eg, Hum = 30%, Wind 5 m/s

02:50-04:00 - Testing new version of scriptor calibration scripts. Did 0.5" zJ and HK200, and 1.5" HK200 (lights off/on worked fine)

04:51 - Open, doing pointing check and collimation

05:09 - N30 camera was in place - had to move back to N3.75

05:10 - resent preset - DIMM 1".24, Avg Guider 1".1

05:21 - UI and Aladdin froze

05:38 - Resending preset - the grating remained in after reset

05:41 - Re-loaded the script and starting again

06:13 - Aladdin in debug mode - waiting to start spectroscopy as we are transiting now

06:20 Move to AO, mostly clear outside

06:23 Steve checks pointing and than did a ACE-preset to A45_R8_0045, Doug checks and put in NCPA, OK

  • Image quality across FOV in JHKs, P_gam, Fe II, Br_gam

Script NGC6819_mod2.xml

06:37 NGC 6819, seeing ~0.6-0.7 on guider, Doug puts in NCPA,

Br_gam, 0252-0256, images slightly elliptical (4.4 vs 4.6pix FWHM ~ 0.067" ) but diffration rings nicely visible

06:53 loop opens, need to restart CCD#39 controller

07:03 Preset again, loop closed
Br_gam, 0257 (w/o NCPA), 0258-0265, last 3 are skies), about 0.065" FWHM,
Fe II, 0266-0273, still about 0.065" FWHM
P_gam, 0274-281 , even in J-range well below 0.07" FWM! Doug tweaks NCPA a bit while doing skies
J, 0282-0285, loop opened, need to resent preset, 0286 crap, 0287 last J on source, 0288-0290 skies
check what images 0291-0293 are

08:13 Gnagnagna...need to resent preset, ima 0294 crap

H, 0295-302, 0.06" FWHM, guider reports 0.7"
Ks, 303-0310, 0.06" FWHM, guider reports 0.65"

Alexander deployes a new SW-version with some more bugs removed

  • Image quality as function of (large) distances in FeII and H-filter

Script NGC6819_wide_mod1.xml

09:50 preset to NGC 6819, 0311

FeII, 0312-0316...Dougs has applied NCPA, but he images did not show this
...typo in Dougs script---

FeII again, ...0318-0322, shell ripped at 0323, Steve took care of it
preset again, 0324 crap, Doug put in NCPA, resumed -offset -60,0 images 0325-331
Image quality at 100" distance still 0.15" or better
H, 0332-0333, FLAO is complaining that offset is out of range of stages. Not true, since RS-star
is at center of detector (but in FLAO-SW?). Preset again, loop closed,
H, 0334-346

  • PSF_test

ES1959+65_PSF_mod1.xml

10:35 Preset to 1ES 1959+650, seeing on guider about 0.7", RS = 15.3 (BIN=4), Dougs tries to start NCPA
Ks, 0347-366

11:17 2nd preset failed - Error due to lack of rotator angle

11:25 - edited file manually, 2nd preset now working. Had to restart scheduler

11:50 Steve's closing up

End of AO-Com#4

-- %USERSIG{JochenHeidt - 2015-06-01}%
Topic revision: r5 - 01 Jun 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback